OpenStudy (anonymous):

Solve the system x + y = 6 and x - y = 2. What is the value of x?

OpenStudy (anonymous):

x=6-y so6-y-y=2 6-2y=2 2y=4 so y=2

OpenStudy (jackellyn):

like @ritvikvij said and then you plug in y in one of the equations So x =6-2=4